OpenStudy (anonymous):

correct my answers? will fan and medal. questions and my answers in comments

OpenStudy (anonymous):

1. Which set of x- and y- values represents a function? A) X, 1, 1.5, 2, 1.5. Y, 0, 2, 4, 3. ~ B) X, 0, 1, 1, 2. Y, 1.8, 3.6, 3.9, 5.4. ~ C) X, 1, 2, 1, 2, Y, 6,3,4, 2. ~ D) X, 1, 3, 5, 7. Y, 1, 11, 11, 20. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ My answer: C ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ 2. In football, a field goal is worth three points. which fuction rule relates the number of field goals, f, and the number of points, p? A) f = 3p ~ B)p = 3f ~ C) p = 3f ~ D) p = 3 +f ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ My answer:D ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ 3. Which function rule is represented by the table? X, 0, 1, 2, 3 Y, 3, 5, 7, 9. A)y = x + 3 ~ B)y = X + 6 ~ C)y = 2x + 3 ~ D)y = 4x + 1 ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ My answer: B ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ 4. Which fuction rule is represented by the table? x, 0, 1, 2, 3. y, 0, -7, -14, -21. A) y = 7x ~ B) y = -7x ~ C) y = x - 8 ~ D) y = 7x - 14 ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ My answer: A ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ 5. What are the values of the fuction y = 3x - 4 for x = 0, 1, 2, and 3? A) -4, -1, 2, 5. ~ B) 0, 3, 6, 9. ~ C) -4, 2, 5, 8. ~ D) 0, 1, 5, 9. ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~ My answer: B ~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

OpenStudy (anonymous):

1. is right 2.C 3.B 4.D 5.B

OpenStudy (anonymous):

okay, thank you very much ^^